40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12546     Entry time: Tue Oct 11 00:43:50 2016     In reply to: 12476     Reply to this: 12549
Author: ericq 
Type: Update 
Category: SUS 
Subject: PRM LR problematic again 

Tonight, and during last week's locking, we noticed something intermittently kicking the PRM. I've determined that PRM's LR OSEM is problematic again. The signal is coming in and out, which kicks the OSEM damping loops. I've had the watchdog tripped for a little bit, and here's the last ten minutes of the free swinging OSEM signal:

Here's the hour trend of the PRM OSEMS over the last 7 days a plot of just LR since the fix on the 9th of September.

It looks like it started misbehaving again on the evening of the 5th, which was right when we were trying to lock... Did we somehow jostle the suspension hard enough to knock the foil cap back into a bad spot?

ELOG V3.1.3-